Im using the pythagorean theorem, I need to find the lenghth of the of the third side: leg 1=10, leg 2= x and the hypotenuse is 30.
Found 2 solutions by richwmiller, edjones:
Answer by richwmiller(17219)

10^2+x^2=30^2

Answer by edjones(8007)

10^2+b^2=30^2
100+b^2=900
b^2=800
b=sqrt(400*2)
b=20sqrt(2)
.
Ed